OpenStudy (anonymous):

Please help! Thank you :) What is –5 • 8 • (–7)? A. –280 B. –240 C. 240 D. 280

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

d

OpenStudy (anonymous):

(- 5 * 8) = -40 * (-7) = 280

OpenStudy (anonymous):

When you multiply by two negatives, you get a positive

OpenStudy (anonymous):

-5x8=-40*-7 =280, as rule negative times a negative is positive.

OpenStudy (anonymous):

oOhh okay i see. Could one of you help me with this other question ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

What is the value of the expression |x + y| when x = –7 and y = 18? A. –25 B. –11 C. 11 D. 25

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Plug in -7 for x and 18 for y and take the absolute value

OpenStudy (anonymous):

how would i do that?

OpenStudy (anonymous):

It's C.

OpenStudy (anonymous):

11

OpenStudy (anonymous):

Negative - a positive is positive.

OpenStudy (anonymous):

so after you do the math, you get 11 and because it is not a negative, you don't have to worry about making the change to the absolute value of a number

OpenStudy (anonymous):

For example, -11 + 3 = -8 Because the negative number has more value then the positive.. Therefore the sum is negative. :)

OpenStudy (mathstudent55):

@xjakester Be careful with your statements. You wrote "Negative - a positive is positive." This is not true.